OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -2x-12/-2

OpenStudy (anonymous):

The distributive property says that if we have an abstract equation such as (a^2+ab), we can remove the a's out, where a is any variable or number. So we could rewrite this as: a(a+b). For your equation, we have: (-2x - 12)/-2 2 is a factor of all involved numbers. So we have: [-2(x+6)]/-2 Now we can simply eliminate the -2 above and below and are left with: x+6 We can double check our work by taking -2(x+6)/-2 and multiplying the -2 through. (-2x+(-12))/-2 = (-2x - 12)/-2 Hope this helps. http://www.tutorsean.net

jimthompson5910 (jim_thompson5910):

\[\frac{-2x-12}{-2}\] \[-\frac{1}{2}(-2x-12)\] \[-\frac{1}{2}(-2x)-\frac{1}{2}(-12)\] \[\frac{-(-2x)}{2}+\frac{-(-12)}{2}\] \[\frac{2x}{2}+\frac{12}{2}\] \[x+6\] So \[\frac{-2x-12}{-2}=x+6\]